OpenStudy (anonymous):

I have the answercan someone please check me.. Find the derivative of the function j(x)=square root x^10-x^8

OpenStudy (anonymous):

\[j(x)=\sqrt{x ^{10}-x ^{8}}\]

OpenStudy (anonymous):

What answer did you get?

OpenStudy (anonymous):

I got \[\frac{ 5x^9-4x^7 }{ (x ^{10}-x^8)^{\frac{ 3 }{ 2 }}}\]

OpenStudy (anonymous):

Is that the right answer

OpenStudy (anonymous):

I got x=0.

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

your answer was close. you need to use the formula if x^10-x^8 = u, then j(x)=u^1/2 dj(x)/dx = (1/2)*(u^-1/2)*du and du = 10*x^9-8*x^7 so you can simplify the expression: (10*x^9-8*x^7)/(2*sqrt(x^10-x^8))

OpenStudy (anonymous):

your answer was only slightly different from the correct answer. the 3/2 should be a 1/2